Enter an equation in point-slope form for the line (7,8) and (8,2) is on the line
Found 2 solutions by Alex.33, Alan3354:
Answer by Alex.33(110)

k=(8-2)/(7-8)=-6
(y-2)=-6(x-8)
Hope it helps!

Answer by Alan3354(69443)

Enter an equation in point-slope form for the line (7,8) and (8,2) is on the line
---------
|x   y   1|
|7   8   1| = 0
|8   2   1|

============
x*(8-2) - y*(7-8) + 1*(14-64) = 0
6x + y = 50